# Parabolic Equations with Singular Coefficients and Boundary Data: Analysis and Numerical Simulations

### v1

@@ Abstract.
We investigate linear parabolic equations in divergence form with singular coefficients and non-smooth boundary data. When the diffusion, drift, or potential terms, as well as the initial or boundary conditions, are distributions rather than functions, classical and weak solution concepts become inadequate due to the ill-posedness of products involving distributions. To overcome this, we introduce a framework of very weak solutions based on regularization techniques and the theory of moderate nets. Existence of very weak solutions is established under minimal regularity assumptions. We further prove consistency with classical solutions when the data are smooth and demonstrate uniqueness via negligibility arguments. Finally, we present numerical computations that illustrate the robustness of the very weak solution framework in handling highly singular inputs, including delta-type potentials and distributional boundary traces.
@@
parabolic equation, divergence form, shift term, energy methods, Galerkin approximation, singular coefficients, regularisation, very weak solution.
@@ 1. Introduction
Here, $a(t,x)$ is the (possibly singular) diffusion coefficient associated with the divergence term, $b(t,x)$ represents the drift field, $q(x)$ is a spatial potential, $f(t,x)$ is an external source term, and $u_{0}(x)$, $g_{0}(t)$, and $g_{1}(t)$ are the prescribed initial and boundary data, respectively.
The well-posedness of such parabolic problems with regular coefficients and boundary data is a classical subject in the theory of partial differential equations (see, for instance, [6, 7, 8, 13]). However, when the diffusion coefficient $a$, the drift term $b$, the potential $q$, the source term $f$, the initial data $u_{0}$, or the boundary data $g_{0},g_{1}$ are allowed to be non-regular (e.g., involving Dirac delta functions or other distributions), the standard functional frameworks break down. In such cases, the product of distributions becomes undefined, and well-posedness in the classical or even weak sense fails to hold.
From an analytical standpoint, the study of PDEs with distributional coefficients lies at the frontier of modern functional analysis and the theory of generalized functions. The classical theory, built upon Sobolev spaces and variational methods, becomes inadequate in the presence of distributional multiplicative terms due to Schwartz’s impossibility theorem [28].
Similar challenges arise in hyperbolic equations. To address them, the notion of very weak solutions was introduced in [19], enabling analysis of second-order equations with singular coefficients. This approach has since been developed in numerous works addressing various types of singularities and equations (see [1, 2, 3, 4, 5, 9, 10, 11, 12, 17, 18, 20, 22, 23, 29]). Recent works on elliptic, hyperbolic and parabolic equations for the Sturm-Liouville operator with singular coefficients [21, 24, 25, 26] have successfully employed separation of variables to construct very weak solutions.
In particular, the parabolic case presents additional challenges due to the time evolution aspect, which complicates the treatment of singularities in a dynamic context. The extension of the very weak solution framework to such settings is both analytically rich and methodologically nontrivial. In our approach, we employ regularisation techniques to treat singular coefficients and data, and we derive uniform a priori estimates to control the behaviour of solution sequences. Moreover, we establish consistency with classical solutions in the smooth-data regime, ensuring the robustness and completeness of our methodology.
@@
The purpose of this work is to extend the framework of very weak solutions to the linear parabolic equation (1)–(1.3) in the presence of both spatial and time-dependent singularities. Our analysis accommodates singularities in the diffusion coefficient $a$, drift term $b$, potential $q$, source term $f$, initial data $u_{0}$, and boundary data $g_{0},g_{1}$, by regularising the problematic terms and constructing solution sequences with uniform estimates. This strategy allows us to prove the existence, uniqueness, and consistency of very weak solutions with classical solutions when the data are smooth.
In a recent work by the authors [27], the analysis of the Klein–Gordon equation with singular coefficients was carried out under inhomogeneous Dirichlet boundary conditions. The analytical techniques developed in that study are applicable to the present problem, with appropriate modifications.
In this work, we extend the analytical framework to parabolic partial differential equations with singular coefficients and inhomogeneous Dirichlet boundary conditions in the one-dimensional spatial setting. Our formulation accommodates singularities in all components of the model—including diffusion, drift, potential, source terms, and boundary data—through a regularization strategy. We establish well-posedness in the very weak sense under this framework. Additionally, we present numerical simulations that illustrate the behaviour of the solutions. Our methodology builds upon and adapts analytical techniques previously developed for the Klein–Gordon equation with singularities [27].
The primary novelty of this work lies in the analysis of a parabolic equation with singular coefficients and a source term, under inhomogeneous Dirichlet boundary conditions — a setting which, to the best of our knowledge, has not yet been investigated in the literature. This study provides a robust theoretical framework for future analytical and numerical developments in this direction.
@@ 2. Homogeneous Dirichlet boundary condition
A function $u\in L^{2}(0,T;H^{1}_{0}(0,1))\cap C([0,T];L^{2}(0,1))$ is said to be a weak solution to the problem (2.1), if the following identity holds for all test functions $\varphi\in H^{1}_{0}(0,1)$ and almost every $t\in[0,T]$:
$\frac{d}{dt}\int_{0}^{1}u(t,x)\varphi(x)\,dx$ $+\int_{0}^{1}a(t,x)\,\partial_{x}u(t,x)\,\partial_{x}\varphi(x)\,dx+\int_{0}^{1}b(t,x)\,\partial_{x}u(t,x)\,\varphi(x)\,dx$
$+\int_{0}^{1}q(x)\,u(t,x)\,\varphi(x)\,dx=\int_{0}^{1}f(t,x)\,\varphi(x)\,dx.$ (2.2)
@@ 2.1. Galerkin Approximation
with initial data $d_{m}^{k}(0)=(u_{0},w_{k})$. This is a linear system of ODEs for the vector $\mathbf{d}_{m}(t)=(d_{m}^{1}(t),\ldots,d_{m}^{m}(t))^{\top}$. Since all coefficients are measurable and bounded on $[0,T]$, standard ODE theory ensures the existence and uniqueness of a solution $\mathbf{d}_{m}\in C^{1}([0,T];\mathbb{R}^{m})$, from which the function $u_{m}$ follows via (2.4). ∎
@@ 2.2. Energy Estimates
$a(t,x)\geq\alpha>0,\quad a,\,\partial_{x}b\in L^{\infty}([0,T];L^{\infty}(0,1)),\quad q\in L^{\infty}(0,1).$ (2.8)
@@
$\|u_{m}\|_{L^{\infty}([0,T];L^{2}(0,1))}+\sqrt{\alpha}\|\partial_{x}u_{m}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\sqrt{q}u_{m}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|\partial_{t}u_{m}\|_{L^{2}([0,T];H^{-1}(0,1))}\leq C\exp\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}$
@@
$\qquad\times\left(\|u_{0}\|_{L^{2}(0,1)}+\|f\|_{L^{2}([0,T];L^{2}(0,1))}\right),$ (2.9)
where the constant $C>0$ depends only on the final time $T$ and the spatial domain $(0,1)$, and is independent of $m$, the coefficients $a,b,q$, and the data $u_{0}$ and $f$.
Proof.
We test the weak form (2.6) with $u_{m}$. This yields
$Re(\partial_{t}u_{m},u_{m})+Re(a\,\partial_{x}u_{m},\partial_{x}u_{m})+Re(b\,\partial_{x}u_{m},u_{m})+Re(qu_{m},u_{m})=Re(f,u_{m}).$
The first term is
$Re(\partial_{t}u_{m},u_{m})=\frac{1}{2}\frac{d}{dt}\|u_{m}\|^{2}_{L^{2}(0,1)}.$
Using $a\geq\alpha>0$, the second term satisfies
$Re(a\,\partial_{x}u_{m},\partial_{x}u_{m})\geq\alpha\|\partial_{x}u_{m}\|^{2}_{L^{2}(0,1)}.$
For the third term, integration by parts gives
$Re(b\,\partial_{x}u_{m},u_{m})=-\frac{1}{2}(\partial_{x}b,|u_{m}|^{2}),$
so
$|(b\,\partial_{x}u_{m},u_{m})|\leq\frac{1}{2}\|\partial_{x}b\|_{L^{\infty}}\|u_{m}\|^{2}_{L^{2}(0,1)}.$
The potential term is
$Re(qu_{m},u_{m})=Re(\sqrt{q}u_{m},\sqrt{q}u_{m})=\|\sqrt{q}u_{m}\|^{2}_{L^{2}(0,1)}.$
And the right-hand side is bounded by
@@
Putting all together, we obtain
$\frac{d}{dt}\|u_{m}\|^{2}_{L^{2}(0,1)}+2\alpha\|\partial_{x}u_{m}\|^{2}_{L^{2}(0,1)}+2\|\sqrt{q}u_{m}\|^{2}_{L^{2}(0,1)}\leq\|f\|^{2}_{L^{2}(0,1)}+C_{1}\|u_{m}\|^{2}_{L^{2}(0,1)},$
where $C_{1}:=1+\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}$.
Letting $\eta(t):=\|u_{m}(t)\|_{L^{2}(0,1)}^{2}$, this gives:
$\eta^{\prime}(t)\leq C_{1}\eta(t)+\|f(t)\|_{L^{2}(0,1)}^{2}.$
Applying Grönwall’s inequality, we obtain:
$\|u_{m}(t)\|_{L^{2}(0,1)}^{2}\leq e^{C_{1}t}\left(\|u_{0}\|_{L^{2}(0,1)}^{2}+\int_{0}^{t}\|f(s)\|_{L^{2}(0,1)}^{2}\,ds\right).$
Taking the supremum over $t\in[0,T]$, we deduce:
[…]
where $C$ depends only on $T$.
$\sqrt{\alpha}\|\partial_{x}u_{m}\|_{L^{2}(0,T;L^{2}(0,1))}+\|\sqrt{q}u_{m}\|_{L^{2}(0,T;L^{2}(0,1))}$
$\qquad\leq C\exp{\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}}\left(\|u_{0}\|_{L^{2}(0,1)}+\|f\|_{L^{2}([0,T];L^{2}(0,1))}\right),$ (2.10)
@@
$Re\langle\partial_{t}u_{m}(t),\phi\rangle$ $=Re(f(t),\phi)-Re(a(t)\,\partial_{x}u_{m}(t),\partial_{x}\phi)$
$-Re(b(t)\,\partial_{x}u_{m}(t),\phi)-Re(q\,u_{m}(t),\phi).$
@@
Applying (2.2) for $u_{m}$ and $\partial_{x}u_{m}$ by integrating with respect to $t$ over $[0,T]$, we conclude
$\|\partial_{t}u_{m}\|_{L^{2}(0,T;H^{-1}(0,1))}$ $\leq C\cdot\exp{\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}}$
[…]
@@
Let $T>0$, and consider the interval $(0,1)\subset\mathbb{R}$. Then, the problem (2.1) admits a unique weak solution
$u\in L^{2}([0,T];H^{1}_{0}(0,1))\,\,\,\text{with}\,\,\,\partial_{t}u\in L^{2}([0,T];H^{-1}(0,1)),\,\,\,\sqrt{q}u\in L^{2}([0,T];L^{2}(0,1)),$
@@
Let $\{u_{m}\}_{m=1}^{\infty}$ be the Galerkin approximations constructed in (2.4). The uniform energy estimate (2.2) ensures that the following sequences are bounded independently of $m$:
$\{u_{m}\}\subset L^{2}([0,T];H^{1}_{0}(0,1)),\quad\{\partial_{t}u_{m}\}\subset L^{2}([0,T];H^{-1}(0,1)),$
$\{\sqrt{q}u_{m}\}\subset L^{2}([0,T];L^{2}(0,1)),\quad\{u_{m}\}\subset L^{\infty}([0,T];L^{2}(0,1)).$
By Banach–Alaoglu and the Aubin–Lions compactness lemma (see [15, Theorem 8, p. 289]), there exists a subsequence $u_{m_{l}}$ and a limit function $u$ such that:
$\left\{\begin{aligned} u_{m_{l}}&\rightharpoonup u&&\text{in }L^{2}([0,T];H^{1}_{0}(0,1)),\\ \sqrt{q}u_{m_{l}}&\rightharpoonup\sqrt{q}u&&\text{in }L^{2}([0,T];L^{2}(0,1)),\\ \partial_{t}u_{m_{l}}&\rightharpoonup\partial_{t}u&&\text{in }L^{2}([0,T];H^{-1}(0,1)),\\ u_{m_{l}}&\to u&&\text{in }L^{2}([0,T];L^{2}(0,1)).\end{aligned}\right$. (2.12)
Using these convergence properties and the weak continuity of all terms in the variational formulation, we can pass to the limit in the Galerkin formulation. Hence, $u$ satisfies the weak form of (2.1).
Uniqueness. Suppose $u\in L^{2}([0,T];H^{1}_{0}(0,1))$ with $\partial_{t}u\in L^{2}([0,T];H^{-1}(0,1))$, and that $u$ is a weak solution corresponding to zero initial data and zero forcing term, i.e., $u_{0}=0$ and $f=0$. Then, for a.e. $t\in[0,T]$, testing the weak formulation with $u(t)$ yields:
$Re\langle\partial_{t}u(t),u(t)\rangle+Re(a\,\partial_{x}u(t),\partial_{x}u(t))+Re(b\,\partial_{x}u(t),u(t))+Re(qu(t),u(t))=0.$
Using the identity $Re\langle\partial_{t}u(t),u(t)\rangle=\frac{1}{2}\frac{d}{dt}\|u(t)\|^{2}_{L^{2}(0,1)}$, and integrating by parts in the convective term:
$Re(b\,\partial_{x}u,u)=-\frac{1}{2}(\partial_{x}b\,,|u|^{2})\leq\frac{1}{2}\|\partial_{x}b\|_{L^{\infty}}\|u\|^{2}_{L^{2}(0,1)}.$
Together with the assumption $a\geq\alpha>0$ and the positivity of $q$, we obtain the differential inequality:
$\frac{1}{2}\frac{d}{dt}\|u(t)\|^{2}_{L^{2}(0,1)}+\alpha\|\partial_{x}u(t)\|^{2}_{L^{2}(0,1)}+\|\sqrt{q}u(t)\|^{2}_{L^{2}(0,1)}\leq\frac{1}{2}\|\partial_{x}b\|_{L^{\infty}}\|u(t)\|^{2}_{L^{2}(0,1)}.$
Applying Grönwall’s inequality and using $u(0)=0$, we conclude that $u(t)\equiv 0$ for all $t\in[0,T]$, proving uniqueness. ∎
Proposition 2.4.
Let $u\in L^{2}([0,T];H^{1}_{0}(0,1))\cap C([0,T];L^{2}(0,1))$ be the weak solution to the parabolic problem (2.1), and assume the coefficients satisfy:
$a,\,\partial_{t}a,\,\partial_{x}a,\,b,\,\partial_{x}b\in L^{\infty}([0,T];L^{\infty}(0,1)),\quad a(t,x)\geq\alpha>0,\quad q\in L^{\infty}(0,1).$
Then, the weak solution to the parabolic equation (2.1) we have the following estimates:
$\Big\{\alpha^{1/2}\|\partial_{x}u\|_{L^{2}([0,T];L^{2}(0,1))};\left.\;\|u\|_{L^{\infty}([0,T];L^{2}(0,1))};\;\|\sqrt{q}u\|_{L^{2}([0,T];L^{2}(0,1))}\right\}$
$\qquad\leq C\exp\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}\left(\|u_{0}\|_{L^{2}(0,1)}+\|f\|_{L^{2}([0,T];L^{2}(0,1))}\right),$ (2.13)
$\|\partial_{t}u\|_{L^{2}([0,T];H^{-1}(0,1))}\leq C\exp\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}$

### v2

@@ Abstract.
We investigate linear parabolic equations in divergence form with singular coefficients and non-smooth initial–boundary data. When the diffusion, drift, or potential terms, as well as the source term and boundary conditions, are distributions rather than functions, classical and weak solution concepts become inadequate, since products involving distributions are not well defined in general. To address this difficulty, we introduce a framework of very weak solutions based on regularisation procedures and the theory of moderate nets. Under minimal regularity assumptions, we establish existence of very weak solutions and prove uniqueness via negligibility arguments. Moreover, in the smooth-data regime, we show consistency with classical weak solutions. Finally, we present numerical experiments illustrating the effectiveness of the proposed framework for highly singular inputs, including delta-type potentials and distributional boundary traces.
@@
parabolic equation, divergence form, drift term, energy methods, Galerkin approximation, singular coefficients, regularisation, very weak solution.
@@ 1. Introduction
Here $a(t,x)$ denotes the diffusion coefficient, $b(t,x)$ the drift field, $q(x)$ a potential term, and $f(t,x)$ an external source. The functions $u_{0}$, $g_{0}$, and $g_{1}$ prescribe the initial and boundary data.
The well-posedness of such parabolic problems with regular coefficients and boundary data is a classical subject in the theory of partial differential equations (see, for instance, [9, 10, 11, 16]). However, when the diffusion coefficient $a$, the drift term $b$, the potential $q$, the source term $f$, the initial data $u_{0}$, or the boundary data $g_{0},g_{1}$ are allowed to be distributional (e.g., involving Dirac delta functions or other distributions), the standard functional frameworks break down. In such cases, the product of distributions becomes undefined, and well-posedness in the classical or even weak sense fails to hold.
From an analytical standpoint, PDEs with distributional coefficients pose substantial challenges within the theory of generalized functions. Classical approaches based on Sobolev spaces and variational methods become inadequate, since products of distributions are not well defined in general, as shown by Schwartz’s impossibility theorem [32].
Similar challenges arise in hyperbolic equations. To address them, the notion of very weak solutions was introduced in [22], enabling analysis of second-order equations with singular coefficients. This approach has since been developed in numerous works addressing various types of singularities and equations (see [1, 2, 3, 4, 5, 6, 7, 8, 12, 13, 14, 15, 20, 21, 24, 26, 27, 33]). Recent works on elliptic, hyperbolic and parabolic equations for the Sturm-Liouville operator with singular coefficients [25, 28, 29, 30] have successfully employed separation of variables to construct very weak solutions.
The parabolic case presents additional difficulties, since singularities interact with the time evolution and uniform estimates must be established already at the level of the regularised problems. Extending the very weak solution framework to this setting therefore requires additional analytical tools. In our approach, we regularise the coefficients and data and derive uniform a priori bounds to control the resulting solution families. Moreover, we prove consistency with classical weak solutions in the smooth-data regime, which confirms the validity of the proposed methodology.
@@
The aim of this work is to extend the very weak solution framework to the parabolic problem (1)–(1.3) in the presence of singular coefficients and inhomogeneous Dirichlet boundary data. We allow distributional diffusion and drift coefficients, singular potentials and source terms, as well as non-smooth initial and boundary conditions. By regularising the coefficients and data and establishing uniform a priori estimates, we prove existence and uniqueness of very weak solutions, together with consistency with classical weak solutions in the smooth regime. Finally, we provide numerical simulations illustrating the behaviour of regularised solutions under delta-type singularities.
The present work continues our recent analysis of the Klein–Gordon equation with singular coefficients under inhomogeneous Dirichlet boundary conditions [31]. Many of the analytical techniques developed there, including the regularisation strategy and the derivation of uniform a priori estimates, provide the foundation for the current study. In contrast to the hyperbolic setting, the parabolic case introduces additional difficulties, since singularities interact with the time evolution and require new arguments.
To the best of our knowledge, well-posedness for parabolic divergence-form equations with distributional coefficients and non-homogeneous Dirichlet boundary data has not previously been established within the very weak solution framework. Thus, the primary novelty of this work lies in the analysis of a parabolic equation with singular diffusion, drift, and potential terms, together with a singular source term, under inhomogeneous Dirichlet boundary conditions.
The paper is organised as follows. In Section 2 we recall the energy estimates for the homogeneous Dirichlet problem. Section 3 contains the existence, uniqueness, and consistency results for very weak solutions under inhomogeneous Dirichlet boundary conditions. Section 4 is devoted to numerical experiments.
@@ 2. Homogeneous Dirichlet boundary condition
Since $q$ may change sign and coercivity may fail, we work in the complex Hilbert spaces $L^{2}(0,1;\mathbb{C})$ and $H_{0}^{1}(0,1;\mathbb{C})$ endowed with the standard sesquilinear $L^{2}$ inner product
$(u,v)_{L^{2}(0,1)}:=\int_{0}^{1}u(x)\,\overline{v(x)}\,dx,\qquad\|u\|_{L^{2}(0,1)}^{2}:=(u,u)_{L^{2}(0,1)}.$
All weak formulations below are understood with this convention.
@@
A function
$u\in L^{2}(0,T;H^{1}_{0}(0,1))\cap L^{\infty}([0,T];L^{2}(0,1)),\qquad\partial_{t}u\in L^{2}(0,T;H^{-1}(0,1)),$
is called a weak solution of (2.1) if $u(0)=u_{0}$ in $L^{2}(0,1)$ and for all $\varphi\in H^{1}_{0}(0,1)$ and for almost every $t\in(0,T)$,
$\langle\partial_{t}u(t),\varphi\rangle_{H^{-1},H^{1}_{0}}$ $+\int_{0}^{1}a(t,x)\,\partial_{x}u(t,x)\,\overline{\partial_{x}\varphi(x)}\,dx+\int_{0}^{1}b(t,x)\,\partial_{x}u(t,x)\,\overline{\varphi(x)}\,dx$
$+\int_{0}^{1}q(x)\,u(t,x)\,\overline{\varphi(x)}\,dx=\int_{0}^{1}f(t,x)\,\overline{\varphi(x)}\,dx.$ (2.2)
@@ 2.1. Galerkin Approximation
with initial data $d_{m}^{k}(0)=(u_{0},w_{k})$. This is a linear system of ODEs for the vector $\mathbf{d}_{m}(t)=(d_{m}^{1}(t),\ldots,d_{m}^{m}(t))^{\top}$. Since all coefficients are measurable and bounded on $[0,T]$, standard ODE theory ensures the existence and uniqueness of a solution $\mathbf{d}_{m}\in C^{1}([0,T];\mathbb{C}^{m})$, from which the function $u_{m}$ follows via (2.4). ∎
@@ 2.2. Energy Estimates
$a(t,x)\geq\alpha>0,\quad a,\,b,\,\partial_{x}b\in L^{\infty}([0,T];L^{\infty}(0,1)),\quad q\in L^{\infty}(0,1),$ (2.8)
and
$q=q_{+}-q_{-},\quad q_{+}\geq 0,\quad q_{-}:=\max\{-q,0\}.$
@@
$\|u_{m}\|_{L^{\infty}([0,T];L^{2}(0,1))}+\sqrt{\alpha}\|\partial_{x}u_{m}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\sqrt{q_{+}}u_{m}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|\partial_{t}u_{m}\|_{L^{2}([0,T];H^{-1}(0,1))}\leq C\cdot\exp{\left\{T\big(\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-}\|_{L^{\infty}(0,1)}\big)\right\}}$
$\times\left(1+\frac{1}{\sqrt{\alpha}}\left(\|a\|_{L^{\infty}([0,T]\times(0,1))}+\|b\|_{L^{\infty}([0,T]\times(0,1))}\right)+\|q\|_{L^{\infty}(0,1)}\right)$
where $C>0$ depends only on $T$ and the spatial domain $(0,1)$, and is independent of $m$. All dependence on the coefficients $a,b,q$ (and on the data $u_{0},f$) is explicitly displayed on the right-hand side of (2.2).
Proof.
We test the weak formulation (2.6) with $u_{m}$. Taking real parts, we obtain
$\operatorname{Re}(\partial_{t}u_{m},u_{m})+\operatorname{Re}(a\,\partial_{x}u_{m},\partial_{x}u_{m})+\operatorname{Re}(b\,\partial_{x}u_{m},u_{m})+\operatorname{Re}(qu_{m},u_{m})=\operatorname{Re}(f,u_{m}).$
The time derivative term satisfies
$\operatorname{Re}(\partial_{t}u_{m},u_{m})=\frac{1}{2}\frac{d}{dt}\|u_{m}\|^{2}_{L^{2}(0,1)}.$
Using the ellipticity condition $a\geq\alpha>0$, the diffusion term yields
$\operatorname{Re}(a\,\partial_{x}u_{m},\partial_{x}u_{m})\geq\alpha\|\partial_{x}u_{m}\|^{2}_{L^{2}(0,1)}.$
For the drift term, integration by parts (using the homogeneous Dirichlet boundary condition) gives
$\operatorname{Re}(b\,\partial_{x}u_{m},u_{m})=-\frac{1}{2}(\partial_{x}b,|u_{m}|^{2}).$
The potential term can be written as
$\operatorname{Re}(qu_{m},u_{m})=(q,|u_{m}|^{2}).$
We decompose the potential as $q=q_{+}-q_{-}$, where $q_{+}\geq 0$ and $q_{-}:=\max\{-q,0\}$. Hence,
$(q,|u_{m}|^{2})=(q_{+},|u_{m}|^{2})-(q_{-},|u_{m}|^{2}).$
Combining the above identities, we arrive at the energy inequality
$\frac{1}{2}\frac{d}{dt}\|u_{m}\|^{2}_{L^{2}(0,1)}$ $+\alpha\|\partial_{x}u_{m}\|^{2}_{L^{2}(0,1)}+(q_{+},|u_{m}|^{2})$
$\leq\operatorname{Re}(f,u_{m})+(q_{-},|u_{m}|^{2})+\frac{1}{2}(\partial_{x}b,|u_{m}|^{2}).$
We estimate the individual terms as follows. First,
$(q_{+},|u_{m}|^{2})=\int_{0}^{1}q_{+}|u_{m}|^{2}\,dx=\|\sqrt{q_{+}}u_{m}\|^{2}_{L^{2}(0,1)}.$
Similarly,
$(q_{-},|u_{m}|^{2})=\|\sqrt{q_{-}}u_{m}\|^{2}_{L^{2}(0,1)}\leq\|q_{-}\|_{L^{\infty}(0,1)}\|u_{m}\|^{2}_{L^{2}(0,1)}.$
For the drift contribution, we have
$\frac{1}{2}(\partial_{x}b,|u_{m}|^{2})\leq\frac{1}{2}\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\|u_{m}\|^{2}_{L^{2}(0,1)}.$
Finally, applying Young’s inequality to the source term yields
$\operatorname{Re}(f,u_{m})\leq\frac{1}{2}\|f\|^{2}_{L^{2}(0,1)}+\frac{1}{2}\|u_{m}\|^{2}_{L^{2}(0,1)}.$
Putting all estimates together, we obtain
$\frac{d}{dt}\|u_{m}\|^{2}_{L^{2}(0,1)}$ $+2\alpha\|\partial_{x}u_{m}\|^{2}_{L^{2}(0,1)}+2\|\sqrt{q_{+}}u_{m}\|^{2}_{L^{2}(0,1)}$
$\leq\|f\|^{2}_{L^{2}(0,1)}+\big(1+\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-}\|_{L^{\infty}(0,1)}\big)\|u_{m}\|^{2}_{L^{2}(0,1)}.$
We introduce the constant
$C_{1}:=1+\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-}\|_{L^{\infty}(0,1)}.$
Setting $\eta(t):=\|u_{m}(t)\|^{2}_{L^{2}(0,1)}$, the above inequality implies
$\eta^{\prime}(t)\leq C_{1}\eta(t)+\|f(t)\|^{2}_{L^{2}(0,1)}.$
By Grönwall’s inequality, for all $t\in[0,T]$,
$\|u_{m}(t)\|^{2}_{L^{2}(0,1)}\leq e^{C_{1}t}\left(\|u_{0}\|^{2}_{L^{2}(0,1)}+\int_{0}^{t}\|f(s)\|^{2}_{L^{2}(0,1)}\,ds\right).$
Taking the supremum over $t\in[0,T]$, we conclude
$\|u_{m}\|^{2}_{L^{\infty}(0,T;L^{2}(0,1))}\leq$ $\;C\exp\!\left\{T\big(\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-}\|_{L^{\infty}(0,1)}\big)\right\}$
@@
where the constant $C$ depends only on $T$.
$\alpha\|\partial_{x}u_{m}\|^{2}_{L^{2}([0,T];L^{2}(0,1))}+2\|\sqrt{q_{+}}u_{m}\|^{2}_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad\leq C\exp{\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-}\|_{L^{\infty}(0,1)}\right\}}$
@@
$\operatorname{Re}\langle\partial_{t}u_{m}(t),\phi\rangle$ $=\operatorname{Re}(f(t),\phi)-\operatorname{Re}(a(t)\,\partial_{x}u_{m}(t),\partial_{x}\phi)$
$-\operatorname{Re}(b(t)\,\partial_{x}u_{m}(t),\phi)-\operatorname{Re}(q\,u_{m}(t),\phi).$
@@
We now bound the potential contribution using the previously obtained estimate $u_{m}\in L^{\infty}([0,T];L^{2}(0,1))$. Indeed,
$\|q\|_{L^{\infty}(0,1)}\|u_{m}(t)\|_{L^{2}(0,1)}\leq\|q\|_{L^{\infty}(0,1)}\|u_{m}\|_{L^{\infty}([0,T];L^{2}(0,1))},$
and therefore
$\big\|\,\|q\|_{L^{\infty}(0,1)}\|u_{m}(\cdot)\|_{L^{2}(0,1)}\big\|_{L^{2}(0,T)}\leq\sqrt{T}\,\|q\|_{L^{\infty}(0,1)}\|u_{m}\|_{L^{\infty}([0,T];L^{2}(0,1))}.$
Using (2.2), (2.2) by taking square root and absorbing all numbers into the constant $C$, we conclude
$\|\partial_{t}u_{m}\|_{L^{2}(0,T;H^{-1}(0,1))}$ $\leq C\cdot\exp{\left\{T\big(\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-}\|_{L^{\infty}(0,1)}\big)\right\}}$
$\times\left(1+\frac{1}{\sqrt{\alpha}}\left(\|a\|_{L^{\infty}([0,T]\times(0,1))}+\|b\|_{L^{\infty}([0,T]\times(0,1))}\right)+\|q\|_{L^{\infty}(0,1)}\right)$
@@
Let $T>0$ and consider the interval $(0,1)\subset\mathbb{R}$. Then the problem (2.1) admits a unique weak solution
$u\in L^{2}([0,T];H_{0}^{1}(0,1))\cap L^{\infty}([0,T];L^{2}(0,1)),\qquad\partial_{t}u\in L^{2}([0,T];H^{-1}(0,1)),$
$\sqrt{q_{+}}\,u\in L^{2}([0,T];L^{2}(0,1)),$
@@
Let $\{u_{m}\}_{m=1}^{\infty}$ be the Galerkin approximations constructed in (2.4). The uniform estimate (2.2) implies that, uniformly in $m$,
  • $u_{m}\ \text{is bounded in }L^{2}(0,T;H_{0}^{1}(0,1))\cap L^{\infty}([0,T];L^{2}(0,1)),$
  • $\partial_{t}u_{m}\ \text{is bounded in }L^{2}([0,T];H^{-1}(0,1)),$
  • $\sqrt{q_{+}}\,u_{m}\ \text{is bounded in }L^{2}([0,T];L^{2}(0,1)).$
By Banach–Alaoglu and the Aubin–Lions lemma (see [18, Theorem 8, p. 289]), there exist a subsequence $(u_{m_{l}})$ and a function $u$ such that